### v1

@@ 4.1 Equivalence of the Estimated ATT Between Conventional and UN–DID Methods with Covariates
Similar to the previous section, we compute the residuals $(post_{t}^{j}-\widehat{post_{t}^{j}})$ for each observation using Equation (47). For an observation in group $j$ during the pre-intervention period, the residual is given by $-\hat{\eta_{0}^{j}}-\hat{\eta_{1}^{j}}X^{j}_{i,t}$. Similarly, for an observation in group $j$ during the post-intervention period, the residual is expressed as $1-\hat{\eta_{0}^{j}}-\hat{\eta_{1}^{j}}X^{j}_{i,t}$. Following this, we move on to derive $\widehat{\gamma_{1}^{j}}$ from Equation (46) using the OLS formula:
$\begin{split}\hat{\gamma}_{1}^{j}=&\frac{\sum_{i}\sum_{t}Y_{i,t}^{j}\left(post^{j}_{t}-\hat{post}^{j}_{t}\right)}{\sum_{i}\sum_{t}\left(post^{j}_{t}-\hat{post}^{j}_{t}\right)^{2}}\\ \Rightarrow\hat{\gamma}_{1}^{j}=&\frac{\sum_{i\in t=0}Y_{i,t}^{j}\left(-\hat{\eta}_{0}^{j}-\hat{\eta_{1}^{j}}X^{j}_{i,t}\right)+\sum_{i\in t=1}Y_{i,t}^{j}\left(1-\hat{\eta}_{0}^{j}-\hat{\eta_{1}^{j}}X^{j}_{i,t}\right)}{\sum_{i\in t=0}\left(-\hat{\eta}_{0}^{j}-\hat{\eta_{1}^{j}}X^{j}_{i,t}\right)^{2}+\sum_{i\in t=1}\left(1-\hat{\eta}_{0}^{j}-\hat{\eta_{1}^{j}}X^{j}_{i,t}\right)^{2}}\\ \Rightarrow\hat{\gamma}_{1}^{j}=&\frac{1-\hat{\eta}_{0}^{j}-\hat{\eta_{1}^{j}}X^{j}_{i,t}}{\sum_{i\in t=0}\left(-\hat{\eta}_{0}^{j}-\hat{\eta_{1}^{j}}X^{j}_{i,t}\right)^{2}+\sum_{i\in t=1}\left(1-\hat{\eta}_{0}^{j}-\hat{\eta_{1}^{j}}X^{j}_{i,t}\right)^{2}}\sum_{i\in t=1}Y_{i,1}\\ &-\frac{\hat{\eta}_{0}^{j}-\hat{\eta_{1}^{j}}X^{j}_{i,t}}{\sum_{i\in t=0}\left(-\hat{\eta}_{0}^{j}-\hat{\eta_{1}^{j}}X^{j}_{i,t}\right)^{2}+\sum_{i\in t=1}\left(1-\hat{\eta}_{0}^{j}-\hat{\eta_{1}^{j}}X^{j}_{i,t}\right)^{2}}\sum_{i\in t=0}Y_{i,0}\end{split}$
$\Rightarrow\hat{\gamma}_{1}^{j}=\frac{1-\hat{\eta}_{0}^{j}-\hat{\eta_{1}^{j}}X^{j}_{i,t}}{F}\sum_{i\in t=1}Y_{i,1}-\frac{\hat{\eta}_{0}^{j}-\hat{\eta_{1}^{j}}X^{j}_{i,t}}{F}\sum_{i\in t=0}Y_{i,0}.$ (48)
@@
$F=\sum_{i\in t=0}\left(-\hat{\eta}_{0}^{j}-\hat{\eta_{1}^{j}}X^{j}_{i,t}\right)^{2}+\sum_{i\in t=1}\left(1-\hat{\eta}_{0}^{j}-\hat{\eta_{1}^{j}}X^{j}_{i,t}\right)^{2}.$
Following the proof in Appendix A, we can show that $\frac{\left(1-\hat{\alpha}_{0}-\hat{\alpha}_{1}-\hat{\alpha}_{2}-\hat{\alpha}_{3}x\right)}{G^{\prime}}$ and $\frac{\left(\hat{\alpha}_{0}+\hat{\alpha}_{2}+\hat{\alpha}_{3}x\right)}{G^{\prime}}$ are equivalent reciprocals of the number of observations in the treated group in the post-intervention period and the pre-intervention period, respectively, $X_{i,t}=x$ (provided the sample size is large enough). Assuming that the law of large numbers hold, $\hat{\gamma^{j}_{1}}$ in Equation (48) is the sample analogue of the following:
$\hat{\gamma}_{1}^{j}=\left[E[Y^{j}_{i,1}|X^{j}_{i,t}]-E[Y^{j}_{i,0}|X^{j}_{i,0}]\right].$ (49)
From Equation (49), we can take the difference between $\hat{\gamma^{T}_{1}}$ and $\hat{\gamma^{C}_{1}}$ from the treated and untreated regressions, respectively. Equation (50) shows that the difference is the sample analogue of the ATT shown in Equation (29):
$\begin{split}\hat{\gamma}_{1}^{T}-\hat{\gamma}_{1}^{C}=\left[E[Y^{T}_{i,1}|X^{T}_{i,t}]-E[Y^{T}_{i,0}|X^{T}_{i,0}]\right]-\left[E[Y^{C}_{i,1}|X^{C}_{i,t}]-E[Y^{C}_{i,0}|X^{C}_{i,0}]\right]\end{split}$
@@
It becomes evident when comparing the two approaches that both methods are capable of estimating the ATT with common slope parameters of the covariates between silos. Now, we will demonstrate that the ATT estimates derived from conventional and UN–DID regressions are equal with time-invariant covariates.
In the context of time-invariant covariates, the regression coefficients for $X_{i}$ in equations (36) and (46) are observed to be 0. This is a consequence of the conditional independence assumption, indicating that $X_{i}$ is independent of $D_{i}$ ($X_{i}\perp D_{i}$). Additionally, because $X_{i}$ is a time-invariant covariate, it is independent of $P_{t}$ as well ($X_{i}\perp P_{t}$). As a result, both $\hat{\alpha_{3}}$ and $\eta^{j}_{1}$ are equal to 0. Therefore, $\hat{\alpha_{3}}=\eta^{j}_{1}=0$. From this result, $\hat{\beta_{3}}$ in Equation (40) can be written as
$\begin{split}\hat{\beta}_{3}=&\frac{\sum_{i\in t=1,T}Y_{i,1}^{T}\left(1-\hat{\alpha}_{0}-\hat{\alpha}_{1}-\hat{\alpha}_{2}\right)}{G}-\frac{\sum_{i\in t=0,T}Y_{i,0}^{T}\left(\hat{\alpha}_{0}+\hat{\alpha}_{2}\right)}{G}\\ &-\frac{\sum_{i\in t=1,C}Y_{i,1}^{C}\left(\hat{\alpha}_{0}+\hat{\alpha}_{1}\right)}{G}+\frac{\sum_{i\in t=0,C}Y_{i,0}^{C}\left(-\hat{\alpha_{0}}\right)}{G}.\end{split}$ (51)
This is the same as the ATT in the absence of covariates for the conventional regression, shown in Equation (14). For the UN–DID regression, $\gamma_{1}^{j}$ in Equation (48) can be rewritten as
$\hat{\gamma}_{1}^{j}=\frac{1-\hat{\eta}_{0}^{j}}{B}\sum_{i\in t=1}Y_{i,1}-\frac{\hat{\eta}_{0}^{j}}{B}\sum_{i\in t=0}Y_{i,0}.$ (52)
This is the same as the ATT in the absence of covariates for the UN–DID regression, shown in Equation (22). Therefore, the ATT estimates from both methods are the same.
The exact numerical equivalence is not maintained when time-varying covariates are present. As noted earlier, both $\hat{\alpha_{3}}$ in Equation (36) and $\hat{\eta_{1}^{j}}$ from Equation (46) are not equal to the true effect of $X_{i,t}$ because the law of large numbers does not hold with small sample size. However, as sample sizes increase, the estimates converge due to the law of large numbers. Thus, the equivalence between conventional and UN–DID estimates holds primarily when dealing with large sample sizes.
@@ 4.2 Monte Carlo Design with Time-Invariant Covariates
In this section, we continue to use the same synthetic dataset as in the preceding section. However, we introduce a modification to our data-generating process, where the outcome variable is now influenced by a time-invariant covariate, denoted by $X_{i}$. In our simulations, the time-invariant covariate is represented by a binary variable indicating gender. Specifically, it assumes a value of 1 if the individual is female, and 0 otherwise. Similar to the previous section, we present two cases. In the first case, we set the true treatment effect at 0. In the second case, we set the true treatment effect at 0.1. Additionally, we assign a common slope parameter of 0.5 to the covariate in both the treated and the control silos.
After generating the data, we estimate the ATT using both the conventional and UN–DID regressions to evaluate their properties with time-invariant covariates. This is repeated 1,000 times for each case, following a methodology similar to the one employed in the previous section. The data-generating process for the two cases are as follows:
@@ 4.3 Results: Time-Invariant Covariate
In subsection 4.1, we showed that both the conventional and the UN–DID methods can recover numerically equivalent values of the ATT when implemented on a poolable dataset. However, the proofs are limited to the estimation of the ATT and do not show the equivalence of the standard errors between the two methods. Through the Monte Carlo Simulation study, we access the equivalence of the estimated standard errors between the conventional and the UN–DID methods using a scatter plot, shown in Figure 5. Panels (a) and (b) depicts the relationship between the conventional standard error (y-axis) and the UN–DID standard errors (x-axis) for Case 3 and Case 4, respectively. Upon reviewing the scatter plots, we note that the standard errors lie along the $45^{\circ}$ line but do not align perfectly with it. This indicates that the standard errors are equivalent but not numerically equal. The same results hold for unequal sample sizes (figure not shown).
@@
Figure 5: Time-Invariant Covariate with Equal Sample Sizes: Scatter Plots
@@
We hypothesize that the standard errors are not numerically equivalent due to small sample sizes. To verify that the two methods converge to the true standard error as sample size increases, we repeat the simulation with sample sizes of 1,000, 2,000, 4,000, 8,000, 10,000, and 50,000 while maintaining equal sample sizes in both silos. We then compute the MSE for both the ATTs and their associated standard errors to look for evidence of convergence. The results are shown in Figure 6. The MSEs on the y-axis are plotted against the sample sizes on the x-axis. The MSE for both the ATT and the standard errors between the UN–DID and the true ATT is illustrated by the red line in Figure 6. On the same axis, we plot the MSE for both ATTs and their standard errors when comparing the conventional method to the true ATT (illustrated by the blue line in Figure 6). Finally, we also plot the MSE of both ATT and standard errors between the conventional and UN–DID methods (represented by the green line in Figure 6).
@@
Figure 6 depicts the MSE of ATTs for Case 3 in Panel (a) and for Case 4 in Panel (b). Panels (c) and (d) showcase the MSE of standard errors for Case 3 and Case 4, respectively. We observe that both ATT and standard error MSEs decrease with larger sample sizes, confirming that both the conventional and UN–DID methods converge to the true values of both the ATT and the standard errors as sample size increases. Notably, the UN–DID standard errors consistently exhibit lower MSE than the conventional estimator, suggesting more reliable inference using the UN–DID method, especially with smaller sample sizes. The graph also illustrates a decreasing MSE between the conventional and UN–DID methods with increasing sample size, indicating convergence between the two estimators. A similar convergence is observed in a case with unequal sample sizes, as shown in Figure 7.
@@ 4.4 Monte Carlo Design with Time-Varying Covariates
In this section, we try to assess the performance of the UN–DID estimator with time-varying covariates. The Monte Carlo simulation design closely resembles the simulation of the previous subsection with a single time-invariant covariate. However, we replace the time-invariant covariate with a time-varying covariate, denoted by $X_{it}$. In our simulations, the time-varying covariate we used is age, which is unaffected by treatment even though it changes with time. Similar to the previous subsection, we analyze two cases where the true effect of the treatment is 0 and 0.1, respectively. A common slope parameter of 0.5 is assigned to the covariate in both the treated and the control silos. The DGPs for the two cases are as follows:
@@ 4.5 Results: Time-Varying Covariate
In subsection 4.1, we have shown that both the conventional and UN–DID methods effectively identify the ATT in the presence of time-varying covariates with common slope parameters. However, it is worth noting that, while the estimates of ATT are close, they are not exactly numerically equal. However, with large sample sizes, the estimates get closer to the true value due to the law of large numbers. Therefore, the equivalence between the conventional and the UN–DID estimates holds only when we have a large enough sample size.
To assess the equivalence of ATT estimates and their associated standard errors, we created scatter plots, Figure 5. In Panel (a), we compare the conventional ATT (y-axis) with the UN–DID ATT (x-axis) for Case 5. Panel (b) shows a similar comparison for Case 6. Additionally, panels (c) and (d) illustrate the relationship between the conventional standard error (y-axis) and the UN–DID standard errors (x-axis) for cases 5 and 6, respectively. Upon inspecting the scatter plots, we observe that the points representing ATT estimates and their associated standard errors align along the $45^{\circ}$ line. However, they are not perfectly aligned, suggesting the estimates of ATT are not numerically equal but are equivalent. This pattern is consistent across both cases. We also find a similar trends when dealing with equal sample sizes (figures not shown).
@@
Similar to the previous subsection, we assess the unbiasedness of the UN–DID estimator using kernel densities, illustrated in Figure 9. The kernel densities reveal that the distribution of ATTs is centered around the true ATT value, indicating unbiasedness with time-varying covariates (common). Compared with earlier figures, there are now some differences between the conventional and UN–DID ATTs.
Additionally, in Figure 10, we observe a decrease in MSE for both the conventional and UN–DID methods as sample size increases. This suggests that the ATT and the standard error between the two methods converge as sample size increases. The same patterns are observed for unequal sample sizes between silos, as shown in Figure 11. Interestingly, the standard errors seem to converge faster to the true values with the UN–DID estimator than they do with the conventional estimator when the silos are of unequal size.
@@
4.6 Additional Types of Covariates
These simulations have considered only cases where we have a balanced panel. In practice, many researchers include covariates when they have access to repeated cross-sectional data. With repeated cross-sectional data, when binary or fixed effect-type coefficients are added, the equivalence of the estimated ATTs between UN–DID and conventional DID remains unchanged. The standard errors can be somewhat different. In this case, we recommend estimating
$\equiv\;Y^{T}_{i,t}=\gamma_{0}^{T}+\gamma_{1}^{T}post_{t}^{T}+\gamma_{2}^{i}X^{T}_{i,t}+\nu^{T}_{i,t}$ (53)
$\equiv\;Y^{C}_{i,t}=\gamma_{0}^{C}+\gamma_{1}^{C}post_{t}^{C}+\gamma_{2}^{i}X^{C}_{i,t}+\nu^{C}_{i,t}$ (54)
in place of the models without a constant. In this case, the ATT is just
$\widehat{ATT}=\widehat{\gamma^{T}_{1}}-\widehat{\gamma^{C}_{1}}.$ (55)
The equivalence between this specification and the regressions without a constant has been shown in Section 3.1. Additionally, the standard error is
$\widehat{SE}(\widehat{ATT})=\sqrt{\widehat{SE}(\widehat{\gamma^{T}_{1}})^{2}+\widehat{SE}(\widehat{\gamma^{C}_{1}})^{2}}.$ (56)
5 UN–DID with Staggered Adoption
In this section, we relax Assumption 5 to accommodate a scenario where multiple groups adopt treatment at different points in time, creating a staggered adoption framework. We also impose an additional restriction called absorptive state, which implies that, once a unit is treated, it remains treated for the rest of the period of the study. Similar to Callaway and Sant’Anna (2021), our proposed method will decompose the dataset into several $2\times 2$ group–time cells.
We will introduce some additional notation following Callaway and Sant’Anna (2021). Let $G_{i,g}=1$ if unit $i$ is treated at time $g$, and 0 otherwise. This will be referred to as the treatment start-time dummies. Following Callaway and Sant’Anna (2021), we group together each individual into $g\in G$ cohorts based on the treatment start-time dummies. For a particular group $g$, any period where $t<g$ are called pre-treatment periods. Under the no-anticipation assumption, the ATT(g,t) for pre-treatment periods are 0.
Each group–time cell contains a group that is currently treated and a group that has not yet been treated. The pre-intervention period in each $2\times 2$ cell is defined as the time period right before the treated group is treated ($g-1$). For each $2\times 2$ group–time cell, we will run the following two regressions:

### v2

@@ 4.2 Equivalence of the Estimated ATT Between Conventional DID and UN-DID Methods with Covariates
Similar to the previous section, we compute the residuals $(post_{t}^{j}-\widehat{post}_{t}^{j})$ for each observation using Equation (45). For an observation in group $j$ during the pre-intervention period, the residual is given by $-\widehat{\eta}_{0}^{j}-\widehat{\eta}_{1}^{j}X^{j}_{i,t}$. Similarly, for an observation in group $j$ during the post-intervention period, the residual is expressed as $1-\widehat{\eta}_{0}^{j}-\widehat{\eta}_{1}^{j}X^{j}_{i,t}$. Following this, we move on to derive $\widehat{\gamma}_{1}^{j}$ from Equation (44) using the OLS formula:
$\begin{split}\widehat{\gamma}_{1}^{j}=&\frac{\sum_{i}\sum_{t}Y_{i,t}^{j}\left(post^{j}_{t}-\widehat{post}^{j}_{t}\right)}{\sum_{i}\sum_{t}\left(post^{j}_{t}-\widehat{post}^{j}_{t}\right)^{2}}\\ \Rightarrow\widehat{\gamma}_{1}^{j}=&\frac{\sum_{i\in t=0}Y_{i,t}^{j}\left(-\widehat{\eta}_{0}^{j}-\widehat{\eta}_{1}^{j}X^{j}_{i,t}\right)+\sum_{i\in t=1}Y_{i,t}^{j}\left(1-\widehat{\eta}_{0}^{j}-\widehat{\eta}_{1}^{j}X^{j}_{i,t}\right)}{\sum_{i\in t=0}\left(-\widehat{\eta}_{0}^{j}-\widehat{\eta}_{1}^{j}X^{j}_{i,t}\right)^{2}+\sum_{i\in t=1}\left(1-\widehat{\eta}_{0}^{j}-\widehat{\eta}_{1}^{j}X^{j}_{i,t}\right)^{2}}\\ \Rightarrow\widehat{\gamma}_{1}^{j}=&\frac{1-\widehat{\eta}_{0}^{j}-\widehat{\eta}_{1}^{j}X^{j}_{i,t}}{\sum_{i\in t=0}\left(-\widehat{\eta}_{0}^{j}-\widehat{\eta}_{1}^{j}X^{j}_{i,t}\right)^{2}+\sum_{i\in t=1}\left(1-\widehat{\eta}_{0}^{j}-\widehat{\eta}_{1}^{j}X^{j}_{i,t}\right)^{2}}\sum_{i\in t=1}Y_{i,1}\\ &-\frac{\widehat{\eta}_{0}^{j}-\widehat{\eta}_{1}^{j}X^{j}_{i,t}}{\sum_{i\in t=0}\left(-\widehat{\eta}_{0}^{j}-\widehat{\eta}_{1}^{j}X^{j}_{i,t}\right)^{2}+\sum_{i\in t=1}\left(1-\widehat{\eta}_{0}^{j}-\widehat{\eta}_{1}^{j}X^{j}_{i,t}\right)^{2}}\sum_{i\in t=0}Y_{i,0}\end{split}$
$\Rightarrow\widehat{\gamma}_{1}^{j}=\frac{1-\widehat{\eta}_{0}^{j}-\widehat{\eta}_{1}^{j}X^{j}_{i,t}}{F}\sum_{i\in t=1}Y_{i,1}-\frac{\widehat{\eta}_{0}^{j}-\widehat{\eta}_{1}^{j}X^{j}_{i,t}}{F}\sum_{i\in t=0}Y_{i,0}.$ (46)
@@
$F=\sum_{i\in t=0}\left(-\widehat{\eta}_{0}^{j}-\widehat{\eta}_{1}^{j}X^{j}_{i,t}\right)^{2}+\sum_{i\in t=1}\left(1-\widehat{\eta}_{0}^{j}-\widehat{\eta}_{1}^{j}X^{j}_{i,t}\right)^{2}.$
Following the proof in Appendix A, we can show that $\frac{\left(1-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{2}-\widehat{\alpha}_{3}x\right)}{G^{\prime}}$ and $\frac{\left(\widehat{\alpha}_{0}+\widehat{\alpha}_{2}+\widehat{\alpha}_{3}x\right)}{G^{\prime}}$ are equivalent reciprocals of the number of observations in the treated group in the post-intervention period and the pre-intervention period, respectively, $X_{i,t}=x$ (provided the sample size is large enough). Assuming that the law of large numbers hold, $\widehat{\gamma}^{j}_{1}$ in Equation (46) is the sample analogue of the following:
$\widehat{\gamma}_{1}^{j}=\left[E[Y^{j}_{i,1}|X^{j}_{i,t}]-E[Y^{j}_{i,0}|X^{j}_{i,0}]\right].$ (47)
From Equation (47), we can take the difference between $\widehat{\gamma}^{T}_{1}$ and $\widehat{\gamma}^{C}_{1}$ from the treated and untreated regressions, respectively. Equation (48) shows that the difference is the sample analogue of the ATT shown in Equation (29):
$\begin{split}\widehat{\gamma}_{1}^{T}-\widehat{\gamma}_{1}^{C}=\left[E[Y^{T}_{i,1}|X^{T}_{i,t}]-E[Y^{T}_{i,0}|X^{T}_{i,0}]\right]-\left[E[Y^{C}_{i,1}|X^{C}_{i,t}]-E[Y^{C}_{i,0}|X^{C}_{i,0}]\right]\end{split}$
@@
It becomes evident when comparing UN-DID and conventional DID that both methods are capable of estimating the ATT with common slope parameters of the covariates across groups. Now, we will demonstrate that the ATT estimates derived from conventional and UN-DID regressions are equal with time-invariant covariates.
In the context of time-invariant covariates, the regression coefficients for $X_{i}$ in equations (34) and (44) are observed to be 0. This is a consequence of the conditional independence assumption, indicating that $X_{i}$ is independent of $D_{i}$ ($X_{i}\perp D_{i}$). Additionally, because $X_{i}$ is a time-invariant covariate, it is independent of $P_{t}$ as well ($X_{i}\perp P_{t}$). As a result, both $\widehat{\alpha}_{3}$ and $\widehat{\eta}^{j}_{1}$ are equal to 0. Therefore, $\widehat{\alpha_{3}}=\widehat{\eta}^{j}_{1}=0$. From this result, $\hat{\beta_{3}}$ in Equation (38) can be written as
$\begin{split}\widehat{\beta}_{3}=&\frac{\sum_{i\in t=1,T}Y_{i,1}^{T}\left(1-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{2}\right)}{G}-\frac{\sum_{i\in t=0,T}Y_{i,0}^{T}\left(\widehat{\alpha}_{0}+\widehat{\alpha}_{2}\right)}{G}\\ &-\frac{\sum_{i\in t=1,C}Y_{i,1}^{C}\left(\widehat{\alpha}_{0}+\widehat{\alpha}_{1}\right)}{G}+\frac{\sum_{i\in t=0,C}Y_{i,0}^{C}\left(-\widehat{\alpha}_{0}\right)}{G}.\end{split}$ (49)
This is the same as the ATT in the absence of covariates for the conventional DID regression, shown in Equation (14). For the UN-DID regression, $\gamma_{1}^{j}$ in Equation (46) can be rewritten as
$\widehat{\gamma}_{1}^{j}=\frac{1-\widehat{\eta}_{0}^{j}}{B}\sum_{i\in t=1}Y_{i,1}-\frac{\widehat{\eta}_{0}^{j}}{B}\sum_{i\in t=0}Y_{i,0}.$ (50)
This is the same as the ATT in the absence of covariates for the UN-DID regression, shown in Equation (22). Therefore, the ATT estimates from both methods are the same.
The exact numerical equivalence is not maintained when time-varying covariates are present. As noted earlier, both $\widehat{\alpha}_{3}$ in Equation (34) and $\widehat{\eta}_{1}^{j}$ from Equation (44) are not equal to the true effect of $X_{i,t}$ because the law of large numbers does not hold with small sample size. However, as sample sizes increase, the estimates converge due to the law of large numbers. Thus, the equivalence between conventional DID and UN-DID estimates with time-varying covariates holds primarily with large sample sizes.
@@ 4.3 Monte Carlo Design with Time-Invariant Covariates
In this section, we continue to use the same synthetic dataset described in Section 3.2. We introduce a modification to our data-generating process, where the outcome is now influenced by a time-invariant covariate, denoted by $X_{i}$. In our simulations, the time-invariant covariate is represented by a binary variable indicating gender. Specifically, it assumes a value of 1 if the individual is female, and 0 otherwise. Similar to the previous section, we present two cases. In the first case, we set the true treatment effect at 0. In the second case, we set the true treatment effect at 0.1. Additionally, we assign a common slope parameter of 0.5 to the covariate in both the treated and the control silos.
After generating the data, we estimate the ATT using both the conventional and UN-DID regressions to evaluate their properties with time-invariant covariates. This is repeated 1,000 times for each case, following the methodology employed above. The data-generating processes for the two cases are as follows:
@@ 4.4 Results: Time-Invariant Covariate
In subsection 4.2, we showed that both the conventional DID and the UN-DID methods can recover numerically equivalent values of the ATT when implemented on a poolable dataset. However, the proofs are limited to the estimation of the ATT and do not show the equivalence of the standard errors between the two methods. Through the Monte Carlo Simulation study, we assess the equivalence of the estimated standard errors between the conventional and the UN-DID methods using a scatter plot, shown in Figure 5. Panels (a) and (b) depict the relationships between the conventional DID standard error (y-axis) and the UN-DID standard error (x-axis) for Case 3 and Case 4 with unequal sample sizes, respectively. Reviewing the scatter plots, we note that the standard errors lie along the $45^{\circ}$ line but do not align perfectly with it. This indicates that the standard errors are equivalent but not numerically equal. The same results hold for equal sample sizes (figure not shown).
@@
Figure 5: Time-Invariant Covariate with Unequal Sample Sizes: Scatter Plots
We assess the unbiasedness of the UN-DID estimator with unequal sample sizes and time-invariant covariates using kernel densities, illustrated in Figure 6. The kernel densities show that the distribution of ATTs is centered on the true ATT value, implying unbiasedness with time-invariant covariates. As in Figure 2, the blue line is not visible as it is perfectly underneath the red line (except at the tails), demonstrating numerical equivalence.
Figure 6: Time-Invariant Covariates with Unequal Sample Sizes: Kernel Density
@@
We hypothesize that the standard errors are not numerically equivalent due to small sample sizes. To verify that the two methods converge to the true standard error as sample size increases, we repeat the simulation with sample sizes of 1,000, 2,000, 4,000, 8,000, 10,000, and 50,000 while maintaining equal sample sizes in both silos. We then compute the MSE for both the ATTs and their associated standard errors for each pair, looking for evidence of convergence in the results shown in Figure 7. The MSEs on the y-axis are plotted against the sample sizes on the x-axis. The MSEs between the UN-DID and the true ATT are illustrated by the red lines in Figure 7. On the same axis, we plot the MSEs comparing the conventional DID to the true ATT (the blue lines in Figure 7) and the MSEs between the conventional DID and UN-DID methods (the green lines in Figure 7).
@@
Figure 7 depicts the MSE of ATTs for Case 3 in Panel (a) and for Case 4 in Panel (b). Panels (c) and (d) showcase the MSE of standard errors for Case 3 and Case 4, respectively. We observe that both ATT and standard error MSEs decrease with larger sample sizes, confirming that both the conventional and UN-DID methods converge to the true values of both the ATT and the standard errors as sample size increases. The blue lines (which illustrate the MSEs for the conventional DID vs. true values of ATT and SE) are not visible, as they lie perfectly underneath the red line (MSE for the UN-DID vs. true values of ATT and SE). This also shows the equivalence between the conventional DID and UN-DID ATT and SEs with time-invariant covariates. Figure 7 also illustrates a near zero MSE between the conventional DID and UN-DID methods, indicating the equivalence between the two estimators. The same pattern is observed in a case with unequal sample sizes, as shown in Figure 8.
@@ 4.5 Monte Carlo Design with Time-Varying Covariates
In this section, we assess the performance of the UN-DID estimator with time-varying covariates. The Monte Carlo simulation design closely resembles the simulation of the previous subsection, however we replace the time-invariant covariate with a time-varying covariate, denoted by $X_{it}$. In our simulations, the time-varying covariate is age, which is unaffected by treatment even though it changes with time. Similar to the previous subsection, we analyze two cases where the true effect of the treatment is 0 and 0.1, respectively. A common slope parameter of 0.5 is assigned to the covariate in both the treated and the control silos. The DGPs for the two cases are as follows:
@@ 4.6 Results: Time-Varying Covariate
In subsection 4.2, we showed that both the conventional and UN-DID methods effectively identify the ATT in the presence of time-varying covariates with common slope parameters. The ATT estimates are close (not exactly numerically equal), and converge to the true value in larger samples due to the law of large numbers. Therefore, the equivalence between the conventional DID and the UN-DID estimates holds in larger samples.
To assess the equivalence of ATT estimates and their associated standard errors, we created scatter plots (Figure 9). In Panel (a), we compare the conventional DID ATT (y-axis) with the UN-DID ATT (x-axis) for Case 5 with unequal sample sizes. Panel (b) shows a similar comparison for Case 6. Additionally, panels (c) and (d) illustrate the relationship between the conventional DID standard errors (y-axis) and the UN-DID standard errors (x-axis) for cases 5 and 6, respectively. Upon inspecting the scatter plots, we observe that the points representing ATT estimates and their associated standard errors align along the $45^{\circ}$ line. However, they are not perfectly aligned, suggesting the ATT estimates are not numerically equal but are equivalent. This pattern is consistent across both cases. We also find a similar pattern with equal sample sizes (figures not shown).
@@
We assess the unbiasedness of the UN-DID estimator with unequal sample sizes and time-varying covariates using kernel densities, illustrated in Figure 10. The kernel densities reveal that the distribution of ATTs is centered around the true ATT value, indicating UN-DID’s unbiasedness with time-varying covariates. Compared with Figures 2 and 6, there are now some small differences between the conventional DID and UN-DID ATTs.
In Figure 11, we observe convergence toward zero in the MSEs for both the conventional DID and UN-DID methods compared to the true value as sample size increases with equal sample sizes. The same patterns are observed for unequal sample sizes between silos (Figure 12).
@@
4.7 Differing Covariates by Silo
One interesting feature of siloed data is that the information may be collected and coded differently according to the processes in place in each jurisdiction. Each silo is primarily interested in analyzing the data in isolation for their own purposes, not in comparisons with other jurisdictions. It is therefore likely that not all variables will be standardized or harmonized across silos. UN-DID is designed for settings where data are unpoolable in the sense that they cannot be combined on one server or in one regression, but where the information they capture is similar enough that fundamental issues of data harmonization are not a problem. It is convenient, therefore, that UN-DID can accomodate settings where data are sufficiently harmonized so that variables capture the same concepts across silos, even if they are coded somewhat differently.
Consider, for instance, a variable like family income. Ideally this variable includes the same types of income (wages, investment earnings, government transfers, etc.) and is continuous and not-censored in both silos. However, it is plausible that both silos record family income as categorical: one using bins in increments of \$10,000, while the other uses increments of \$25,000. This would cause issues for conventional DID, as the categories are non-overlapping across silos without a problematic degree of aggregation. With UN-DID, the silo-specific coefficients can handle the different categories easily, each first-difference can be estimated conditional on the covariate, and the conditional ATT and SEs can also be estimated. While data must be sufficiently harmonized to capture the same underlying concepts and the common causal covariate assumption must hold, UN-DID can more flexibly accomodate differences in datasets across data silos than conventional DID.
5 UN-DID with Staggered Adoption
In this section, we relax Assumption 5 to accommodate a scenario where multiple groups adopt treatment at different points in time, creating a staggered adoption framework. We also impose an additional restriction called absorptive state, which implies that, once a unit is treated, it remains treated for the rest of the period of the study. Similar to Callaway and Sant’Anna (2021), UN-DID with staggered adoption will decompose the dataset into several $2\times 2$ group–time cells. Staggered adoption means that there are at least three silos under analysis. Section 6 gives additional guidance for working with multiple silos.
We introduce some additional notation following Callaway and Sant’Anna (2021). Let $G_{i,g}=1$ if unit $i$ is treated at time $g$, and 0 otherwise. These will be referred to as the treatment start-time dummies. Following Callaway and Sant’Anna (2021), we group together each individual into $g\in G$ cohorts based on the treatment start-time dummies. For a particular group $g$, any period where $t<g$ are called pre-treatment periods. Under the no-anticipation assumption, the ATT(g,t)s for pre-treatment periods are 0.
Each group–time cell contains a group that is currently treated and a group that has not yet been treated. If treated and control observations within this cell are in separate data silos, we can implement UN-DID to estimate the ATT(g,t). The pre-intervention period in each $2\times 2$ cell is defined as the time period right before the treated group is treated ($g-1$). For each $2\times 2$ group–time cell, we will run the following two regressions: